OpenStudy (anonymous):

What are fractions that are equivalent to 1.2 and 2.7 ( those number have the dash on the top like - to say they repeat forever) Thank you :)

OpenStudy (anonymous):

\[\overline{.2}=\frac{2}{9}\]

OpenStudy (anonymous):

\[\overline{.7}=\frac{7}{9}\]

OpenStudy (anonymous):

\[1.2 = \frac{11}{9}\] and \[2.7 = \frac{25}{9}\]

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

Can i get a medal por favor? O.o

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Ty!

OpenStudy (anonymous):

Welcome :)